# Preconditioned Score-based Generative Models

### v2

@@ 3.2 Sampling from the Langevin dynamics
where the initialization $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},I)$, $\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t-1})$ is approximated by $\mathbf{s}_{\theta}(\mathbf{x}_{t-1},\sigma_{i})$ at a proper level of $\sigma_{i}$, $\epsilon_{t}>0$ the step size, and $T$ the iterations. With this process, we transform a Gaussian sample $\mathbf{x}_{0}$ to $\mathbf{x}_{T}$ in the target distribution $p_{\mathbf{x}^{\ast}}$. The Langevin dynamics Eq. (10) works in that as $T\rightarrow+\infty$, the distribution of $\mathbf{x}_{t}$ converges to a steady-state distribution, exactly $p_{\mathbf{x}^{\ast}}$ [25]. Improvements can be made by refining noise scales, iterations and step size, and using the moving average technique [3].
@@ 3.3 Reverse diffusion
where the initial distribution $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},I)$, $\bar{g}_{t}=g_{T-t}$, $\bar{p}_{t}={p}_{T-t}$ is the probabilistic distribution function of $\mathbf{x}_{T-t}$ in the forward process. In this way, the final-steady state $\mathbf{x}_{T}$ of reverse diffusion obeys $p_{\mathbf{x}^{\ast}}$ [50]. Eq. (11) and (12) both are the discretized version of the corresponding stochastic differential equation (SDE). Hence, a numerical solver of SDE can be used to implement the sampling process, as realized in NCSN++ [4]. Specifically, NCSN++ learns to fit the gradient term $\bigtriangledown_{\mathbf{x}}\log p_{T-t}(\mathbf{x})$ through minimizing
$\mathbb{E}_{t}\lambda_{t}\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathbf{x}^{\ast}}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}(t),t)-\bigtriangledown_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],$ (13)
where $t$ is uniformly drawn from $\{1,\ldots,T\}$, and $\lambda_{t}$ is the weight at $t$. During sampling, NCSN++ updates the state in two steps (Alg. 1): (1) predicting with the reverse diffusion (Eq. (12)), (2) correcting with the Langevin dynamics (Eq. (10)). This improves both accuracy and stability. The variance exploding (VE) variant is the best among all NCSN++ models [4].
Algorithm 1 The sampling process of VE NCSN++
@@ 3.4 Limitation analysis
While having superior generating capability over GANs [1], SGMs run much slower. A direct way for acceleration is to reduce the iterations $T$ and increase the step size $\epsilon_{t}$ proportionally (i.e., one step for multiple ones). This however leads to severe degradation (see Fig. 1). This is partly because the sampling process is fundamentally limited in sampling structural data (e.g., images), as elaborated below.
Theorem 1.
@@
As $U$ is orthogonal, $UU^{\mathsf{T}}=I_{d}$. We thus finish the proof for the Langevin dynamics Eq. (10). The proof of the reverse diffusion Eq.(12) is almost the same. ∎
@@
Since any permutation matrix is orthogonal, Eq. (10) ignores the coordinate order of data. This property could be suitable for unstructured data, but not for structured data (e.g., images) where the coordinate order represents the geometry information of individual pixels locally and the semantic structures globally.
3.5 Leveraging preconditioning
Conceptually, the Langevin dynamics Eq. (10) of the sampling process is a special case of Metropolis adjusted Langevin algorithm (MALA) [23, 24, 25] with non-autonomous factor $\epsilon_{t}$. Inspired by the findings that the preconditioning approach can significantly accelerate the MALA [23], we study how to construct a preconditioning operator $M$ to accelerate the sampling phase of SGMs.
Specifically, the original sample space may suffer from an ill-conditioned issue caused by the large variations of the scales in different coordinates. This could lead to oscillation and slow convergence, especially when the step size $\epsilon_{t}$ is not small enough [51]. Preconditioning can make the scales become more similar along all the coordinates [23, 26] using a matrix preconditioner, alleviating the ill-conditioned problem and improving the convergence of the sampling process even under a large step size. This applies to the reverse diffusion (Eq. (12)) similarly.
As shown in the following sections, with the proposed preconditioning, the noise terms $\mathbf{z}_{t}$ become anisotropic (i.e., its covariance matrix is no longer a scalar matrix). As a result, the sampling process of SGMs is varying under the permutation of coordinates, becoming conditional to the coordinate order (Thm. 1).
@@ 4 Theoretical analysis
In this section, we develop the theories for the formulation of our preconditioning operator suitable for the predicting-correcting based sampling process (Alg. 1). As predicting and correcting can be applied in isolation for complete sampling [4], we will inspect them separately.
@@ 4.1 Continuous processes
starting from an initial distribution $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},I)$, with $\mathbf{w}_{t}$ a Wiener process, and $\epsilon_{t}$ the continuous version of the step size. The continuous version for corrupting process (Eq. (11)) is
@@
starting from an initial distribution $\mathbf{x}_{0}\sim p_{\mathbf{x}^{\ast}}$. The continuous version of reverse diffusion (Eq. (12))
@@
starting from an initial distribution $\mathbf{x}_{0}\sim p_{\mathbf{x}^{\ast}}$. Note, with these processes, the distribution of the initial state $\mathbf{x}_{0}$ is different.
4.2 Preserving the steady-state distribution for the Langevin dynamics
To ensure the generation quality, the steady-state distribution of the Langevin dynamics (Eq. (18)) must stick with the target distribution $p_{\mathbf{x}^{\ast}}$; Otherwise, poor performance will be likely resulted. We prove that our preconditioning preserves the steady-state distribution.
We first derive the theory behind the Langevin dynamics on why it can generate samples from $p_{\mathbf{x}^{\ast}}$. According to [27], Eq. (18) is associated with a Fokker-Planck equation
@@ 4.2 Preserving the steady-state distribution for the Langevin dynamics
where $p=p(\mathbf{x},t)$ is the distribution of $\mathbf{x}$ at time $t$. This describes the stochastic dynamics of Eq. (18).
Let $\frac{\partial p}{\partial t}=0$, we obtain the steady-state function
@@
$\bigtriangledown_{\mathbf{x}}\cdot(\bigtriangledown_{\mathbf{x}}\log(p_{\mathbf{x}^{\ast}})p_{\mathbf{x}^{\ast}})=\bigtriangledown_{\mathbf{x}}\cdot(\bigtriangledown_{\mathbf{x}}(p_{\mathbf{x}^{\ast}}))=\Delta_{\mathbf{x}}p^{\ast},$ (23)
which is exactly the R.H.S. of the equation. Hence, the target data distribution $p_{\mathbf{x}^{\ast}}$ is one of the steady-state solutions of Eq. (18). It is proved that this steady-state solution is unique, s.t. $\epsilon_{t}>0$ [52]. As a result, with enough iterations, the state distribution will converge to the target data distribution $p_{\mathbf{x}^{\ast}}$.
@@
where $M$ is an invertible linear operator, $S$ is any skew-symmetric operator, and $M^{\mathrm{T}}$ is the adjoint operator of $M$.
Theorem 2.
@@
Proof.
The invertibility of $M$ preserves the uniquity of the solution of the steady-state equation [27]. Hence we only need to prove that the target data distribution $p_{\mathbf{x}^{\ast}}$ still satisfies the steady-state equation after preconditioning. The Fokker-Planck equation of Eq. (24) is
@@
since $S$ is skew-symmetric. Therefore, the term $\bigtriangledown_{\mathbf{x}}\cdot(S\bigtriangledown_{\mathbf{x}}p_{\mathbf{x}^{\ast}})=0$. Then, the steady-state solution of Eq. (22), i.e., $p_{\mathbf{x}^{\ast}}$, also satisfies the steady-state equation of Eq. (24). As a result, the theorem is proved. ∎
Thm. 2 provides a wide manipulative space for us to design the preconditioning operator $M$ and $S$ while preserving the steady-state distribution of $\mathbf{x}_{T}$ in the Langevin dynamics simultaneously.
4.3 Preserving the final-state distribution for the reverse diffusion
Unlike the Langevin dynamics, the reverse diffusion (Eq. (20)) directly reverses the corrupting process. So its final-steady state $\mathbf{x}_{T}$ obeys $p_{\mathbf{x}^{\ast}}$ naturally [50]. Note the final-steady distribution of Eq. (20) may not be its steady-state distribution. This can be seen if we write out its steady-state equation through Fokker-Planck equation
@@ 4.3 Preserving the final-state distribution for the reverse diffusion
Accordingly, to enable the reverse diffusion generate samples from $p_{\mathbf{x}^{\ast}}$ after preconditioning, we must preserve its final-state distribution $\mathbf{x}_{T}$. To that end, we modify Eq. (20) as
@@
where $M$ is an invertible linear operator. In fact, we have
Theorem 3.
@@
Denote $\tilde{\bar{p}}_{t}(\mathbf{y})$ as the probabilistic distribution function of the $\mathbf{y}_{T-t}$ in the process Eq. (26). Denote $\mathbf{y}_{t}=M^{-1}\mathbf{x}_{t}$, and substituting $\mathbf{y}_{t}$ into Eq. (25), we have
@@
$M^{-T}\bigtriangledown_{\mathbf{x}_{t}}\log\bar{p}_{t}(\mathbf{x}_{t})=\bigtriangledown_{\mathbf{y}_{t}}\log\tilde{\bar{p}}_{t}(\mathbf{y}_{t}).$ (29)
@@
which is exactly the reverse process of Eq. (26[50]. As a result, the final-state distribution of Eq. (27) obeys $\mathbf{y}_{T}\sim p_{\mathbf{y}^{\ast}}$. Reversing $M$ on Eq. (27), the final-state distribution of Eq. (25) thus obeys $\mathbf{x}_{T}\sim p_{\mathbf{x}^{\ast}}$. ∎
Similar as Thm. 2, Thm. 3 provides a wide space for designing the preconditioner $M$ while preserving the final-state distribution of $\mathbf{x}_{T}$ in the reverse diffusion.
@@ 5 Preconditioned diffusion sampling
Thus far, we have presented the generic preconditioning formulation for both Langevin dynamics and reverse diffusion. Next, we focus on image generation tasks by designing proper operators.
@@ 5.1 Image preconditioning strategies
In numerical optimization, accurate preconditioning operators could be obtained by calculating the inverse of the Hessian matrix of $\log p_{\mathbf{x}^{\ast}}(\mathbf{x})$ at each time step. This however is extremely costly for high dimensional tasks (e.g., image generation) [53]. To bypass this obstacle, we adopt a data-centred strategy that leverages the per-coordinate scale information of target data.
For imagery data, we consider two types of coordinate: individual pixels (i.e., the spacial coordinate) and frequency (i.e., Fourier transformed). The former is useful for images with global structure characteristics, as per-pixel statistics gives prior information (i.e., pixel preconditioner). With natural images, the amplitude of low-frequency components is typically much higher than that of high-frequency ones [54]. This structural frequency property could be similarly useful. (i.e., frequency preconditioner). For either case, we construct the linear operator $M$ with the mean value of target training data, used to balance all the coordinates (Fig. 2). We simply set the skew-symmetric operator $S$ (Eq. (24)) as zero and will discuss it later.
Combining the two precondtioners, we form a linear preconditioning operator $M\in\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as $M[\cdot]=M_{f}[M_{p}[\cdot]]$ with $M_{f}$ and $M_{p}$ the preconditioner for frequency and pixel, respectively. We operate $M$ on the noise term $\mathbf{z}_{t}$ and adjust the gradient term to keep the steady-state distribution for the Langevin dynamics Eq. (10) as
$\mathbf{x}_{t}=\mathbf{x}_{t-1}+\frac{\epsilon_{t}^{2}}{2}M_{p}[M_{f}^{2}[M_{p}[\bigtriangledown_{\mathbf{x}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t-1})]]]$
$+\epsilon_{t}M_{f}[M_{p}[\mathbf{z}_{t}]],1\leq t\leq T.$ (30)
@@
$\mathbf{x}_{t}=\mathbf{x}_{t-1}+\bar{g}_{t}^{2}M_{p}[M_{f}^{2}[M_{p}[\bigtriangledown_{\mathbf{x}}\log\bar{p}_{t}(\mathbf{x}_{t-1})]]]$
$+\bar{g}_{t}M_{f}[M_{p}[\mathbf{z}_{t}]],1\leq t\leq T,$ (31)
We called this proposed sampling procedure as preconditioned diffusion sampling (PDS), as summarized in Alg. 2.

### v3

@@ 3.2 Langevin dynamics
where the initialization $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},I)$, $\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t-1})$ is approximated by $\mathbf{s}_{\theta}(\mathbf{x}_{t-1},\sigma_{i})$ at a proper level of $\sigma_{i}$, $\epsilon_{t}>0$ the step size, and $T$ the iterations. With this process, we transform a Gaussian sample $\mathbf{x}_{0}$ to $\mathbf{x}_{T}$ in the target distribution $p_{\mathbf{x}^{\ast}}$. Note that we define our notations for the forward and reverse processes in the manner that the process begins with $\mathbf{x}_{0}$ and ends with $\mathbf{x}_{T}$. The Langevin dynamics Eq. (8) works in that as $T\rightarrow+\infty$, the distribution of $\mathbf{x}_{t}$ converges to a steady-state distribution, exactly $p_{\mathbf{x}^{\ast}}$ (Girolami and Calderhead, 2011).. Improvements can be made by refining noise scales, iterations, and step size, and using the moving average technique (Song and Ermon, 2020).
@@ 3.3 Reverse diffusion
where the initial distribution $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},I)$, $\bar{g}_{t}=g_{T-t}$, $\bar{p}_{t}={p}_{T-t}$ is the probabilistic distribution function of $\mathbf{x}_{T-t}$ in the forward process. In this way, the final-steady state $\mathbf{x}_{T}$ of reverse diffusion obeys $p_{\mathbf{x}^{\ast}}$ (Haussmann and Pardoux, 1986). Eq. (9) and (10) both are the discretized version of the corresponding stochastic differential equation (SDE). Hence, a numerical solver of SDE can be used to implement the sampling process, as realized in NCSN++ (Song et al., 2021b). Specifically, NCSN++ learns to fit the gradient term $\bigtriangledown_{\mathbf{x}}\log p_{T-t}(\mathbf{x})$ through minimizing
$\mathbb{E}_{t}\lambda_{t}\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathbf{x}^{\ast}}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}(t),t)-\bigtriangledown_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],$
where $t$ is uniformly drawn from $\{1,\ldots,T\}$, and $\lambda_{t}$ is the weight at $t$. During sampling, NCSN++ updates the state in two steps (Alg. 1): (1) predicting with the reverse diffusion (Eq. (10)), (2) correcting with the Langevin dynamics (Eq. (8)). The variance exploding (VE) variant is the best among all NCSN++ models (Song et al., 2021b).
Algorithm 1 Sampling process of VE NCSN++
@@ 3.4 Limitation analysis
While having superior generating capability over GANs, SGMs run much slower. A direct way for acceleration is to reduce the iterations $T$ and increase the step size $\epsilon_{t}$ proportionally (i.e., one step for multiple ones). This however leads to severe degradation (see Fig. 1). This is partly because the sampling process is fundamentally limited in sampling structural data (e.g., images), as elaborated below.
Theorem 3.1
@@
As $U$ is orthogonal, $UU^{\mathsf{T}}=I_{d}$. We thus finish the proof for the Langevin dynamics Eq. (8). The proof of the reverse diffusion Eq. (10) is almost the same.
@@
Since any permutation matrix (which operates on a vector by changing the order of its coordinates) is orthogonal, Thm. 3.1 implies that the sampling process remains equivalent regardless of how the coordinates of $\mathbf{x}_{t}$ are reordered. However, for structured data (e.g., images), the order of coordinates encodes both local geometric information of each pixels and global semantic structures. Therefore, the current sampling process in Eq. (8) disregards the structural integrity of the data.
3.5 Preconditioning
Conceptually, the Langevin dynamics Eq. (8) of the sampling process is a special case of Metropolis adjusted Langevin algorithm (MALA) (Roberts and Stramer, 2002; Welling and Teh, 2011; Girolami and Calderhead, 2011) with non-autonomous factor $\epsilon_{t}$. Inspired by the findings that the preconditioning approach can significantly accelerate the MALA (Roberts and Stramer, 2002), we study how to construct a preconditioning operator $M$ to accelerate the sampling phase of SGMs.
Specifically, the original sample space may suffer from an ill-conditioned issue caused by the large variations of the scales in different coordinates. This could lead to oscillation and slow convergence, especially when the step size $\epsilon_{t}$ is not small enough (Nocedal and Wright, 1999). Preconditioning can make the scales become more similar along all the coordinates (Roberts and Stramer, 2002; Li et al., 2016) using a matrix preconditioner, alleviating the ill-conditioned problem and improving the convergence of the sampling process even under a large step size. This applies to the reverse diffusion (Eq. (10)) similarly. As shown in the following sections, with the proposed preconditioning, the noise terms $\mathbf{z}_{t}$ become anisotropic (i.e., its covariance matrix is no longer a scalar matrix). As a result, the sampling process of SGMs varies under the permutation of coordinates, becoming conditional to the coordinate order (Thm. 3.1).
@@ 4 Theoretical analysis
In this section, we develop the theories for the formulation of our preconditioning operator suitable for the sampling process incorporating prediction and correction (Alg. 1). Since the sampling process consists of two types of stochastic processes—prediction (reverse diffusion) and correction (Langevin dynamics)—we must demonstrate that the preconditioning operator preserves the steady-state distribution of each process separately.
@@ 4.1 Continuous processes
starting from an initial distribution $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},I)$, with $\mathbf{w}_{t}$ a Wiener process, and $\epsilon_{t}$ is the continuous version of the step size. The continuous version for corrupting process (Eq. (9)) is
@@
starting from an initial distribution $\mathbf{x}_{0}\sim p_{\mathbf{x}^{\ast}}$. The continuous version of reverse diffusion is(Eq. (10))
@@
starting from an initial distribution $\mathbf{x}_{0}\sim p_{\mathbf{x}^{\ast}}$. Note that these processes result in a different distribution for the initial state $\mathbf{x}_{0}$.
4.2 Preserving the steady-state distribution of the Langevin dynamics
To ensure generation quality, the steady-state distribution of the Langevin dynamics (Eq. (15)) must match the target distribution $p_{\mathbf{x}^{\ast}}$. Failure to satisfy this condition may lead to poor performance. We prove that the proposed preconditioning preserves the steady-state distribution.
We first derive the theory behind the Langevin dynamics on why it can generate samples from $p_{\mathbf{x}^{\ast}}$. According to (Gardiner et al., 1985), Eq. (15) is associated with a Fokker-Planck equation
@@ 4.2 Preserving the steady-state distribution of the Langevin dynamics
where $p=p(\mathbf{x},t)$ is the distribution of $\mathbf{x}$ at time $t$. This describes the stochastic dynamics of Eq. (15). Let $\frac{\partial p}{\partial t}=0$, we obtain the steady-state function
@@
$\bigtriangledown_{\mathbf{x}}\cdot(\bigtriangledown_{\mathbf{x}}\log(p_{\mathbf{x}^{\ast}})p_{\mathbf{x}^{\ast}})=\bigtriangledown_{\mathbf{x}}\cdot(\bigtriangledown_{\mathbf{x}}(p_{\mathbf{x}^{\ast}}))=\Delta_{\mathbf{x}}p^{\ast},$
which is exactly the R.H.S. of the equation. Hence, the target data distribution $p_{\mathbf{x}^{\ast}}$ is one of the steady-state solutions of Eq. (15). It is proved that this steady-state solution is unique, s.t. $\epsilon_{t}>0$ (Risken and Eberly, 1996). As a result, with enough iterations, the state distribution will converge to the target data distribution $p_{\mathbf{x}^{\ast}}$.
@@
where $M$ is an invertible linear operator, $M^{\mathrm{T}}$ is its adjoint operator, and $S$ is a skew-symmetric operator.
Theorem 4.1
@@
Proof
The invertibility of $M$ preserves the uniqueness of the solution of the steady-state equation (Gardiner et al., 1985). Hence, we only need to prove that the target data distribution $p_{\mathbf{x}^{\ast}}$ still satisfies the steady-state equation after preconditioning. The Fokker-Planck equation of Eq. (20) is
@@
since $S$ is skew-symmetric. Therefore, the term $\bigtriangledown_{\mathbf{x}}\cdot(S\bigtriangledown_{\mathbf{x}}p_{\mathbf{x}^{\ast}})=0$. Then, the steady-state solution of Eq. (19), i.e., $p_{\mathbf{x}^{\ast}}$, also satisfies the steady-state equation of Eq. (20). As a result, the theorem is proved.
Thm. 4.1 provides a wide manipulative space for us to design the preconditioning operator $M$ and $S$ while preserving the steady-state distribution of $\mathbf{x}_{T}$ in the Langevin dynamics simultaneously. More precisely, given a trained SGM with the sampling process described by Eq. (15), the sampling process can be modified to Eq. (4.1) without altering the final generated distribution, provided that the step size is sufficiently small.
4.3 Preserving the final-state distribution of the reverse diffusion
Although seemingly similar, the reverse diffusion (Eq. (17)) is different from the Langevin dynamics (Eq. (15)) as $\bar{p}_{t}(\mathbf{x})\neq p_{\mathbf{x}^{\ast}}(\mathbf{x})$ in general. Therefore, we need to reanalyze how to preserve the steady-state distribution of the reverse diffusion process when introducing alteration to the process. As the reverse diffusion (Eq. (17)) directly reverses the corrupting process, its final-steady state $\mathbf{x}_{T}$ obeys $p_{\mathbf{x}^{\ast}}$ naturally (Haussmann and Pardoux, 1986). Note that the final-steady distribution of Eq. (17) may not be its steady-state distribution. This can be seen if we write out its steady-state equation through the Fokker-Planck equation
@@ 4.3 Preserving the final-state distribution of the reverse diffusion
Accordingly, to enable the reverse diffusion to generate samples from $p_{\mathbf{x}^{\ast}}$ after preconditioning, we must preserve its final-state distribution $\mathbf{x}_{T}$. To that end, we modify Eq. (17) as
@@
In fact, we have
Theorem 4.2
@@
Denote $\tilde{\bar{p}}_{t}(\mathbf{y})$ as the probabilistic distribution function of the $\mathbf{y}_{T-t}$ in the process Eq. (22). Denote $\mathbf{y}_{t}=M^{-1}\mathbf{x}_{t}$, and substitute $\mathbf{y}_{t}$ into Eq. (21), we have
@@
$M^{-T}\bigtriangledown_{\mathbf{x}_{t}}\log\bar{p}_{t}(\mathbf{x}_{t})=\bigtriangledown_{\mathbf{y}_{t}}\log\tilde{\bar{p}}_{t}(\mathbf{y}_{t}).$
@@
which is exactly the reverse process of Eq. (22(Haussmann and Pardoux, 1986). As a result, the final-state distribution of Eq. (23) obeys $\mathbf{y}_{T}\sim p_{\mathbf{y}^{\ast}}$. Reversing $M$ on Eq. (23), the final-state distribution of Eq. (21) thus obeys $\mathbf{x}_{T}\sim p_{\mathbf{x}^{\ast}}$.
Similar to Thm. 4.1, Thm. 4.2 provides a wide space for designing the preconditioner $M$ while preserving the final-state distribution of $\mathbf{x}_{T}$ in the reverse diffusion.
@@ 5 Preconditioned diffusion sampling
Thus far, we have presented the generic preconditioning formulation for both Langevin dynamics and reverse diffusion. Before designing the preconditioned diffusion sampling for image generation, we first test its efficiency through a toy example.
5.1 Toy example
We implement Langevin dynamics sampling to generate samples from a two-dimensional Gaussian distribution $\mathcal{N}(\mathbf{0},C)$, where the covariance matrix is $C=\begin{bmatrix}9.0&0.2\\ 0.2&0.5\end{bmatrix}$. The condition number of the covariance matrix is $18.18$, indicating ill-conditioning. As a common preconditioning strategy, we set the preconditioning matrix for this sampling process as $C+I$. As shown in the left and middle of Fig. 3, if we increase the step size, the original Langevin dynamics fail to generate samples of high probability. Whereas the preconditioned Langevin dynamics produces samples with much higher probability with large step size, as shown in the right of Fig. 3.
Figure 3: Sampling on a two-dimensional Gaussian distribution using Langevin dynamics with step size $0.1$ (left), Langevin dynamics with step $1$ (middle) and preconditioned Langevin dynamics with step $1$ (right).
Next, we focus on image generation tasks by designing proper operators. Unlike the toy model, image data are generally high-dimensional. Using the same strategy as the toy model—relying on the data’s covariance matrix—would significantly increase computational costs during sampling, as matrix multiplication becomes computationally expensive in high-dimensional spaces. To overcome this, we leverage the structure of image data and design specialized preconditioners that enable efficient computation.
@@ 5.2 Image preconditioning strategies
In numerical optimization, accurate preconditioning operators is calculated from the inverse of the Hessian matrix of $\log p_{\mathbf{x}^{\ast}}(\mathbf{x})$ at each time step. This however is extremely costly for high dimensional tasks (e.g., image generation) (Wright et al., 1999). To bypass this obstacle, we adopt a data-centred strategy that leverages the coordinate scale information of target data.
For imagery data, we consider two types of coordinates: individual pixels (i.e., the spacial coordinate) and frequency (i.e., Fourier transformed). The former is useful for images with global structure characteristics, as per-pixel statistics gives prior information (i.e., pixel preconditioner). With natural images, the amplitude of low-frequency components is typically much higher than that of high-frequency ones (van der Schaaf and van Hateren, 1996). This structural frequency property could be similarly useful. (i.e., frequency preconditioner). For either case, we construct the linear operator $M$ with the mean value of target training data, used to balance all the coordinates (Fig. 2). We set the skew-symmetric operator $S$ (Eq. (20)) as zero and will discuss it later. Combining the two preconditioners, we form a linear preconditioning operator $M\in\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as $M[\cdot]=M_{f}[M_{p}[\cdot]]$ with $M_{f}$ and $M_{p}$ the preconditioner for frequency and pixel, respectively. We operate $M$ on the noise term $\mathbf{z}_{t}$ and adjust the gradient term to keep the steady-state distribution for the Langevin dynamics Eq. (8) as
[…]
@@
[…]
We call this sampling procedure as preconditioned diffusion sampling (PDS), as summarized in Alg. 2.